QC is a median of triangle PQR. The incircle of triangle PQR intersects this median at points A and B, and at that A lies between Q and B, and QA : AB : BC=1:2:1. Find the ratio of the largest side of triangle PQR to its smallest side.


1
Expert's answer
2021-10-29T04:35:37-0400

If the incircle intersect median of a triangle it has to be a right angle triangle and because its ratio is 1:2:1 the triangle is to be isosceles right angle triangle

therefore the smaller side be 2A

so "2a^2+2a^2=b^2" (here b is hypotenuse )

so "\frac{b}{a}= 2 \sqrt 2"